OpenStudy (anonymous):

f(x)with radical sign 0.5x-5;f(20),f(16),f(12),f(14) f(20)= find domain of the function

OpenStudy (anonymous):

can any one solve this

OpenStudy (mertsj):

\[f(x)=\sqrt{.5x-5}\]

OpenStudy (mertsj):

Is that the function?

OpenStudy (mertsj):

If it is the domain is all the numbers that make .5x-5 positive or 0

OpenStudy (mertsj):

\[\frac{1}{2}x-5\ge0\]

OpenStudy (mertsj):

\[x \ge10\]

OpenStudy (mertsj):

That is the domain.

OpenStudy (mertsj):

\[f(20)=\sqrt{.5(20)-5}=\sqrt{5}\]

OpenStudy (mertsj):

\[f(16)=\sqrt{.5(16)-5}=\sqrt{8-5}=\sqrt{3}\]

OpenStudy (mertsj):

\[f(12)=\sqrt{.5(12)-5}=\sqrt{6-5}=\sqrt{1}=1\]

OpenStudy (mertsj):

\[f(14)=\sqrt{.5(14)-5}=\sqrt{7-5}=\sqrt{2}\]

OpenStudy (anonymous):

type an integer or decimal rounded to the nearest tenth as needed

OpenStudy (anonymous):

f(20)=2.2 is the correct answer

OpenStudy (mertsj):

Do you have a calculator?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Then pick it up and type square root 5 and see if you don't get 2.2

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Now do the others in the same way.

OpenStudy (mertsj):

What is the square root of 3?

OpenStudy (anonymous):

1.7

OpenStudy (mertsj):

What about the square root of 2?

OpenStudy (anonymous):

1.4

OpenStudy (mertsj):

So now you have all the answers.

OpenStudy (anonymous):

so domain of f(x) is all real numbers x where ?

OpenStudy (mertsj):

You know, it would really be nice if, when you ask for help, and someone helps you that you would have the common courtesy to read what they wrote.

OpenStudy (anonymous):

1 am with you